# Quantum field theory treatment of neutrino spin-flavor oscillations in a magnetic field

### v1

@@ Abstract
We study spin-flavor oscillations of neutrinos in a magnetic field within the quantum field theory approach in which neutrinos are virtual particles. Neutrinos are taken to be Majorana particles having a nonzero transition magnetic moment. We derive the dressed propagators of the neutrino mass eigenstates exactly accounting for the magnetic field contribution. The matrix element and the transition probability for spin-flavor oscillations are obtained. The leading term in the transition probability is shown to coincide with the result of the standard quantum mechanical description of neutrino oscillations. We also derive the correction to the transition probability which has the quantum field theory origin.
@@ 1 Introduction
The interaction of neutrinos with external fields can significantly modify the dynamics of neutrino oscillations. From the point of view of the phenomenology, the most important external field, contributing to flavor oscillations, is the electroweak interaction of neutrinos with background matter. It can result in the sizable amplification of the transition probability, known as the Mikheyev–Smirnov–Wolfenstein effect [1, 2], which is the most plausible explanation of the solar neutrino deficit [3].
Neutrinos are supposed to be electrically neutral particles. Nevertheless, these particles can interact with electromagnetic fields owing to nonzero magnetic moments which have purely anomalous origin (see, e.g., Ref. [4]). It should be noted that the electromagnetic properties of Dirac and Majorana neutrinos are completely different. Dirac neutrinos can have both diagonal and transition magnetic moments, whereas Majorana neutrinos can possess only transition ones. Despite the issue of the neutrino nature is still open, neutrinos are likely to be Majorana particles. At least, the smallness of the active neutrino masses is explained more naturally for Majorana neutrinos [5]. The recent advances in the neutrinos electromagnetic properties are reviewed in Ref. [6].
If neutrinos with transition magnetic moments interact with an external magnetic field, they undergo both flavor change and the spin flip. This process is called neutrino spin-flavor oscillations [7]. Neutrino spin-flavor oscillations were studied, e.g., in Ref. [8] to explain the possible anticorrelation between the solar activity and the solar neutrinos flux. Later, spin-flavor oscillations were understood in Ref. [9] to play a subdominant role for the solar neutrinos deficit problem. The description of spin-flavor oscillations of Majorana neutrinos expressed in terms of Weyl spinors was carried out in Ref. [10].
As a rule, neutrino oscillations in vacuum and in external fields are described within the quantum mechanical approach, which gives satisfactory results in the majority of situations. Nonetheless, this formalism has some shortcomings which were outlined in Ref. [11]. Numerous attempts to construct the description of neutrino flavor oscillations based on the quantum field theory (QFT) methods have been undertaken. Some of them are reviewed in Ref. [12].
@@
The approach in Refs. [13, 14, 15] allows one to treat neutrino oscillations in vacuum since it involves the propagators of the neutrino mass eigenstates. In vacuum, these propagators are diagonal in the neutrino types. If a neutrino interacts with an external field, the mass eigenstates propagators become nondiagonal. Finding such nondiagonal propagators in not a trivial task. A particular case of spin-flavor oscillations of Dirac neutrinos, possessing diagonal magnetic moments in the mass basis, was analyzed in frames of QFT in Ref. [18].
@@
In the present work, we apply the results of Ref. [19] for the studies of neutrino spin-flavor oscillations in a magnetic field. Moreover, we consider the situation when the interaction with an external field is essentially nondiagonal in the mass basis. We shall see shortly in Sec. (2) that it is implemented for Majorana neutrinos. Therefore, the application of the results of Ref. [19] is crucial in this case. Note that the alternative formalism, based on relativistic quantum mechanics, for the description of neutrino flavor and spin-flavor oscillations in various external fields was considered in Ref. [20].
This work is organized in the following way. We remind briefly some basic issues of the Majorana neutrinos electrodynamics in Sec. 2. Then, in Sec. 3, we outline the QFT approach and adapt it for neutrino spin-flavor oscillations. We derive the dressed propagators for the neutrino mass eigenstates in a magnetic field by solving the corresponding Dyson equations in Sec. 4. The matrix element and the transition probability for neutrino spin-flavor oscillations are computed in Sec. 5. The results of the QFT and the standard quantum mechanical approaches are compared in Sec. 6. Finally, we conclude in Sec. 7. The propagators of massive Weyl neutrinos in vacuum are derived in Appendix.
@@ 3 Quantum field theory formalism for neutrino oscillations
In this section, we briefly outline the quantum field theory based formalism for neutrino oscillations in which neutrinos are virtual particles [13, 14, 15]. We adapt this method for the description of spin-flavor oscillations.
Following Ref. [13], we assume that the source and the detector of neutrinos, separated by the distance $\mathbf{L}$, consist of heavy nuclei $N$. The production and detection of neutrinos is characterized by the interaction of these nuclei with charged leptons $l_{\alpha}^{-}$ and antileptons $l_{\beta}^{+}$, where $\alpha,\beta=e,\mu,\dotsc$. Thus, neutrino oscillations correspond to the process $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+N^{\prime}$. If the flavor content of charged leptons changes, i.e. $\beta\neq\alpha$, one can speak about neutrino flavor oscillations. If electric charges of leptons also change, e.g., we have an electron in a source and antimuon in a detector, such a process is allowed in vacuum for Majorana neutrinos. It was connected to neutrino-to-antineutrino oscillations in Refs. [13, 23]. However, if we study the transition $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+N^{\prime}$ in an external field which neutrinos interact with, it is more appropriate to call it neutrino spin-flavor oscillations. The considered process is schematically depicted in Fig. 1.
Figure 1: The schematic Feynman diagram for the reaction $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+N^{\prime}$. The neutrino line has two counter-directional arrows since we deal with the Majorana particles. Moreover, the neutrino line is broad since the mass eigenstates interact with a magnetic field [see Eq. (2)]. Note that we take into account the magnetic interaction exactly (see Sec. 4).
@@
$S=-\frac{1}{2}\left(\sqrt{2}G_{\mathrm{int}}\right)^{2}\int\mathrm{d}^{4}x\mathrm{d}^{4}y\Big{\langle}l_{\alpha}^{+},N,N^{\prime}\Big{|}T\left\{j^{\mu}(x)J_{\mu}(x)j^{\nu}(y)J_{\nu}(y)\right\}\Big{|}l_{\beta}^{-},N,N^{\prime}\Big{\rangle},$ (3.1)
@@
$l_{\lambda}(x)=\int\mathrm{d}^{3}p\left[a_{\lambda}(p)u_{\lambda}(p)e^{-\mathrm{i}p_{\lambda}x}+b_{\lambda}^{\dagger}(p)v_{\lambda}(p)e^{\mathrm{i}p_{\lambda}x}\right],$ (3.3)
@@
×$\times\left(\int\mathrm{d}^{4}x\mathrm{d}^{4}ye^{\mathrm{i}p_{\alpha}x-\mathrm{i}p_{\beta}y}J_{\mu}(x)J_{\nu}(y)\big{\langle}0\big{|}T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\big{|}0\big{\rangle}\right)\gamma_{\mathrm{L}}^{\nu}u_{\beta}(p_{\beta}),$ (3.4)
where we account for Eqs. (3.2) and (3.3).
@@
$v_{\alpha}^{\mathrm{T}}(p_{\alpha})(\gamma_{\mathrm{L}}^{0})^{\mathrm{T}}\left\langle 0\left|T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\right|0\right\rangle\gamma_{\mathrm{L}}^{0}u_{\beta}(p_{\beta})\\ =\kappa_{\alpha\mathrm{R}}^{\dagger}(p_{\alpha})\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle\kappa_{\beta\mathrm{L}}(p_{\beta}).$ (3.6)
@@
$\psi_{a}=\left(\begin{array}[]{c}\xi_{a}\\ \eta_{a}\end{array}\right),\quad u_{\alpha,\beta}(p_{\alpha,\beta})=\left(\begin{array}[]{c}\kappa_{\alpha,\beta\mathrm{R}}(p_{\alpha,\beta})\\ \kappa_{\alpha,\beta\mathrm{L}}(p_{\alpha,\beta})\end{array}\right).$ (3.7)
@@
$\mathcal{M}_{\beta\to\bar{\alpha}}=\sum_{ab}U_{\alpha a}^{*}U_{\beta b}^{*}\kappa_{\alpha+}^{\dagger}(p_{\alpha})\left(\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\Sigma_{ab}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\right)\kappa_{\beta-}(p_{\beta}),$ (3.9)
where $E=(E_{\alpha}+E_{\beta})/2$ and $\Sigma_{ab}(q_{0},\mathbf{q})$ are the 4D Fourier images of the nondiagonal propagators,
$\Sigma_{ab}(x-y)=\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle,$ (3.10)
We consider utrarelativistic incoming and outgoing leptons. Therefore, in Eq. (3.9), we replace $\kappa_{\beta\mathrm{L}}$ and $\kappa_{\alpha\mathrm{R}}$ with the helicity states $\kappa_{\beta-}$ and $\kappa_{\alpha+}$.
The probability of neutrino spin-flavor oscillations is $P_{\beta\to\bar{\alpha}}\propto|\mathcal{M}_{\beta\to\bar{\alpha}}|^{2}$. In case of the Majorana neutrinos motion in vacuum, the propagators are diagonal, $\Sigma_{ab}\propto\delta_{ab}$. In this situation, the process $\beta\to\bar{\alpha}$ corresponds to neutrino-to-antineutrino oscillations, as we mentioned earlier. The probability of $\nu_{\alpha}\to\bar{\nu}_{\beta}$ oscillations was found in Refs. [13, 23] to be of the order $\sim(m_{1,2}/E)^{2}$.
If Majorana neutrinos have a transition magnetic moment, as one can see in Eq. (2), the wave equations for $\eta_{1}$ and $\eta_{2}$ are coupled by the magnetic interaction. This interaction induces the transitions between $\eta_{1,2}$ and $\mathrm{i}\sigma_{2}\eta_{2,1}^{*}=\xi_{2,1}$. It means that both diagonal, $\Sigma_{11}$ and $\Sigma_{22}$, and nondiagonal, $\Sigma_{12}$ and $\Sigma_{21}$, propagators in Eq. (3.10) can contribute to the transition probability. In case we study spin-flavor oscillations like $\nu_{e}\to\bar{\nu}_{\mu}$, the matrix element in Eq. (3.9) takes the form,
$\mathcal{M}_{e\to\bar{\mu}}=\kappa_{\mu+}^{\dagger}(p_{\mu})\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}e^{\mathrm{i}\mathbf{qL}}\left[\frac{1}{2}\left(\Sigma_{22}-\Sigma_{11}\right)\sin 2\theta+\cos^{2}\theta\Sigma_{21}-\sin^{2}\theta\Sigma_{12}\right]\kappa_{e-}(p_{e}),$ (3.11)
@@ 4 Propagators of massive Majorana neutrinos in a magnetic field
In this section, we derive both diagonal and nondiagonal propagators of Majorana neutrinos interacting with a magnetic field. For this purpose, we use the method developed in Ref. [19], where the propagators in question are the solutions of the Dyson equations exactly accounting for the interaction with an external field. First, we derive the Dyson equations within the perturbative approach, with all the terms in the perturbation series being summed up. Then, the Dyson equations are solved exactly for ultrarelativistic neutrinos. We deal with two mass eigenstates since we mentioned in Ref. [19] that extension of the formalism for the greater number of neutrinos is ambiguous.
@@
Figure 2: Feynman diagrams corresponding to Eq. (4). 4D Fourier images of the bare propagators $-\mathrm{i}S_{a\mathrm{R,L}}$, given in Eqs. (A.2) and (A.10), are depicted by thin solid lines [24]. The broad solid lines correspond to the dressed propagators $-\mathrm{i}\Sigma_{ab}$. The potential of the magnetic interaction is $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$.
The series in Eq. (4) can be summed up. The result is
@@
We notice the diagonal propagators in Eq. (4) do not contribute to the matrix element since $\mathcal{M}_{e\to\bar{\mu}}$ is averaged in Eq. (3.11) over the leptonic states with opposite helicities. Hence, we should look for $\Sigma_{12}$ and $\Sigma_{21}$ only.
One can see in the expressions for $S_{a\mathrm{L,R}}$ in Eqs. (A.2) and (A.10) that these propagators contain the projectors $(1\mp\bm{\sigma}\hat{p})$ for ultrarelativistic neutrinos. Thus, the inverse operators $S_{a\mathrm{L,R}}^{-1}$, which are present in Eq. (4), do not exist for such neutrinos. However, we keep the ratio $\varsigma_{a}=p/E_{a}$ in soling of Eq. (4) and set $\varsigma_{a}\to 1$ in the final result. To proceed we choose the geometry of the system. Let us take that $\mathbf{L}=L\mathbf{e}_{z}$, where $\mathbf{e}_{z}$ is the unit vector along the $z$-axis. We also assume that the process is considered in the forward scattering approximation, i.e. $\mathbf{p}_{e}\upuparrows\mathbf{p}_{\bar{\mu}}\upuparrows\mathbf{e}_{z}$, which is valid for ultrarelativistic leptons. The magnetic field is taken to be transverse to the charged leptons motion, $\mathbf{B}=B\mathbf{e}_{x}$, where $\mathbf{e}_{x}$ is the unit vector along the $x$-axis.
We provide the detailed calculations for $\Sigma_{12}$ only. Using Eqs. (A.2) and (A.10), the exact solution of the Dyson equation for $\Sigma_{12}$ in Eq. (4) has the form,
@@
×$\times\big{\{}\alpha(\varsigma_{1}-\varsigma_{2})\mu(\hat{q}\mathbf{B})+(\bm{\sigma}\cdot[\mathrm{i}\alpha(\varsigma_{1}+\varsigma_{2})\mu(\hat{q}\times\mathbf{B})-2\alpha\varsigma_{1}\varsigma_{2}\mu(\hat{q}\mathbf{B})\hat{q}$
@@
and $E_{1,2}=\sqrt{q^{2}+m_{1,2}^{2}}$. For ultrarelativistic neutrinos, one has that $\varsigma_{1,2}\to 1$ and Eq. (4) takes the form,

### v2

# Quantum field theory treatment of the neutrino spin-flavor precession in a magnetic field

@@ Abstract
We study the spin-flavor precession of neutrinos in a magnetic field within the quantum field theory approach in which neutrinos are virtual particles. Neutrinos are taken to be Majorana particles having a nonzero transition magnetic moment. We derive the dressed propagators of the neutrino mass eigenstates exactly accounting for the magnetic field contribution. The matrix element and the transition probability for the spin-flavor precession are obtained in the approximation of the forwardly scattered charged leptons. The leading term in the transition probability is shown to coincide with the result of the standard quantum mechanical description of neutrino oscillations. We also discuss the quantum field theory contributions to the neutrino dressed propagators and demonstrate that these contributions result in a small correction to the transition probability. The case of charged leptons with arbitrary energies is considered.
@@ 1 Introduction
The interaction of neutrinos with external fields can significantly modify the dynamics of neutrino oscillations. From the point of view of the phenomenology, the most important external field, contributing to flavor oscillations, is the electroweak interaction of neutrinos with background matter. It can result in the sizable amplification of the transition probability, known as the Mikheyev–Smirnov–Wolfenstein (MSW) effect [1, 2], that is the most plausible explanation of the solar neutrino deficit [3].
Neutrinos are supposed to be electrically neutral particles. Nevertheless, these particles can interact with electromagnetic fields owing to nonzero magnetic moments which have a purely anomalous origin (see, e.g., Ref. [4]). It should be noted that the electromagnetic properties of Dirac and Majorana neutrinos are completely different. Dirac neutrinos can have both diagonal and transition magnetic moments, whereas Majorana neutrinos can possess only transition ones. Despite the issue of the neutrino nature is still open, neutrinos are likely to be Majorana particles. At least, the smallness of the active neutrino masses is explained more naturally for Majorana neutrinos [5]. The recent advances in the studies of neutrinos electromagnetic properties are reviewed in Ref. [6].
If neutrinos with transition magnetic moments interact with an external magnetic field, they undergo both flavor change and the spin flip. This process is called the neutrino spin-flavor precession [7]. The neutrino spin-flavor precession was studied, e.g., in Ref. [8] to explain the possible anticorrelation between the solar activity and the solar neutrinos flux. The spin-flavor precession accounting for the neutrino interaction with background matter was shown in Ref. [9] to exhibit a resonant behavior analogous to the MSW effect. The application of the predicted effect to solar neutrinos was also discussed in Ref. [9]. Later, the spin-flavor precession was understood in Ref. [10] to play a subdominant role for the solar neutrinos deficit problem. The description of the spin-flavor precession of Majorana neutrinos expressed in terms of the Weyl spinors was carried out in Ref. [11].
As a rule, neutrino oscillations in vacuum and in external fields are described within the quantum mechanical approach, which gives satisfactory results in the majority of situations. Nonetheless, this formalism has some shortcomings which were outlined in Ref. [12]. Numerous attempts to construct the description of neutrino flavor oscillations based on the quantum field theory (QFT) methods are reviewed in Ref. [13].
@@
The approach in Refs. [14, 15, 16] allows one to treat neutrino oscillations in vacuum since it involves the propagators of the neutrino mass eigenstates. In vacuum, these propagators are diagonal in the neutrino types. If a neutrino interacts with an external field, the mass eigenstates propagators become nondiagonal since external fields mix different mass eigenstates as a rule. Finding such nondiagonal propagators in not a trivial task. A particular case of the spin-flavor precession of Dirac neutrinos, possessing diagonal magnetic moments in the mass basis, was analyzed in frames of QFT in Ref. [19].
@@
In the present work, we apply the results of Ref. [20] for the studies of the neutrino spin-flavor precession in a magnetic field. Moreover, we consider the situation when the interaction with an external field is essentially nondiagonal in the mass basis. We shall see shortly in Sec. 2 that it is implemented for Majorana neutrinos. Therefore, the application of the results of Ref. [20] is crucial in this case. Note that the alternative formalism, based on the relativistic quantum mechanics, for the description of neutrino flavor oscillations and the spin-flavor precession in various external fields was considered in Ref. [21].
This work is organized in the following way. We remind briefly some basic issues of the Majorana neutrinos electrodynamics in Sec. 2. Then, in Sec. 3, we outline the QFT approach and adapt it for the neutrino spin-flavor precession. We derive the dressed propagators for the neutrino mass eigenstates in a magnetic field by solving the corresponding Dyson equations in Sec. 4. The matrix element and the transition probability for the neutrino spin-flavor precession are computed in Sec. 5. The results of the QFT and the standard quantum mechanical approaches are compared in Sec. 6. Finally, we discuss our results in Sec. 7. The propagators of massive Weyl neutrinos in vacuum are derived in Appendix.
@@ 3 Quantum field theory formalism for neutrino oscillations
In this section, we briefly outline the quantum field theory based formalism for neutrino oscillations in which neutrinos are virtual particles [14, 15, 16]. We adapt this method for the description of the spin-flavor precession.
Following Ref. [14], we assume that the source and the detector of neutrinos, separated by the distance $\mathbf{L}$, consist of heavy nuclei $N$ and $N^{\prime}$. The production and detection of neutrinos is characterized by the interaction of these nuclei with charged leptons $l_{\alpha}^{-}$ and antileptons $l_{\beta}^{+}$, where $\alpha,\beta=e,\mu,\dotsc$. Thus, neutrino oscillations correspond to the process $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$. If the flavor content of charged leptons changes, i.e. $\beta\neq\alpha$, one can speak about neutrino flavor oscillations. If electric charges of leptons also change, e.g., we have an electron in a source and antimuon in a detector, such a process is allowed in vacuum for Majorana neutrinos. It was connected to neutrino-to-antineutrino oscillations in Refs. [14, 24]. However, if we study the transition $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$ in an external field which neutrinos interact with, it is more appropriate to call it the neutrino spin-flavor precession. The considered process is schematically depicted in Fig. 1.
Figure 1: The schematic Feynman diagram for the reaction $l_{\beta}^{-}+N\to\text{neutrinos}\to l_{\alpha}^{+}+\tilde{N}^{\prime}$. The neutrino line has two counter-directional arrows since we deal with the Majorana particles. Moreover, the neutrino line is broad since the mass eigenstates interact with a magnetic field [see Eq. (2)]. Note that we take into account the magnetic interaction exactly (see Sec. 4).
@@
$S=-\frac{1}{2}\left(\sqrt{2}G_{\mathrm{int}}\right)^{2}\int\mathrm{d}^{4}x\mathrm{d}^{4}y\Big{\langle}l_{\alpha}^{+},\tilde{N},\tilde{N}^{\prime}\Big{|}T\left\{j^{\mu}(x)J_{\mu}(x)j^{\nu}(y)J_{\nu}(y)\right\}\Big{|}l_{\beta}^{-},N,N^{\prime}\Big{\rangle},$ (3.1)
@@
$l_{\lambda}(x)=\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3/2}}\left[a_{\lambda}(p)u_{\lambda}(p)e^{-\mathrm{i}p_{\lambda}x}+b_{\lambda}^{\dagger}(p)v_{\lambda}(p)e^{\mathrm{i}p_{\lambda}x}\right],$ (3.3)
@@
$\times\left(\int\mathrm{d}^{4}x\mathrm{d}^{4}ye^{\mathrm{i}p_{\alpha}x-\mathrm{i}p_{\beta}y}J_{\mu}(x)J_{\nu}(y)\big{\langle}0\big{|}T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\big{|}0\big{\rangle}\right)\gamma_{\mathrm{L}}^{\nu}u_{\beta}(p_{\beta}),$ (3.4)
where we account for Eqs. (3.2) and (3.3). In Eq. (3), we use the mean values of the nuclear currents $J_{\mu}(x)=\langle\tilde{N}^{\prime}|J_{\mu}|N^{\prime}\rangle$ and $J_{\nu}(y)=\langle\tilde{N}|J_{\nu}|N\rangle$.
@@
$v_{\alpha}^{\mathrm{T}}(p_{\alpha})(\gamma_{\mathrm{L}}^{0})^{\mathrm{T}}\left\langle 0\left|T\left\{(\bar{\psi}_{a}(x))^{\mathrm{T}}\bar{\psi}_{b}(y)\right\}\right|0\right\rangle\gamma_{\mathrm{L}}^{0}u_{\beta}(p_{\beta})\\ =\kappa_{\bar{\alpha}\mathrm{R}}^{\dagger}(p_{\alpha})\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle\kappa_{\beta\mathrm{L}}(p_{\beta}).$ (3.6)
@@
$\psi_{a}=\left(\begin{array}[]{c}\xi_{a}\\ \eta_{a}\end{array}\right),\quad u_{\alpha,\beta}(p_{\alpha,\beta})=\left(\begin{array}[]{c}\kappa_{\bar{\alpha},\beta\mathrm{R}}(p_{\alpha,\beta})\\ \kappa_{\bar{\alpha},\beta\mathrm{L}}(p_{\alpha,\beta})\end{array}\right).$ (3.7)
@@
$\mathcal{M}_{\beta\to\bar{\alpha}}=\sum_{ab}U_{\alpha a}^{*}U_{\beta b}^{*}\kappa_{\bar{\alpha}\mathrm{R}}^{\dagger}(p_{\alpha})\left(\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}\Sigma_{ab}(E,\mathbf{q})e^{\mathrm{i}\mathbf{qL}}\right)\kappa_{\beta\mathrm{L}}(p_{\beta}).$ (3.9)
Here, $E=(E_{\alpha}+E_{\beta})/2$ and $\Sigma_{ab}(q_{0},\mathbf{q})$ are the 4D Fourier images of the nondiagonal propagators,
$\Sigma_{ab}(x-y)=\left\langle 0\left|T\left\{\xi_{a}(x)\eta_{b}^{\dagger}(y)\right\}\right|0\right\rangle.$ (3.10)
If we consider utrarelativistic incoming and outgoing leptons, then, in Eq. (3.9), we should replace $\kappa_{\beta\mathrm{L}}$ and $\kappa_{\bar{\alpha}\mathrm{R}}$ with the helicity states $\kappa_{\beta-}$ and $\kappa_{\bar{\alpha}+}$.
The probability of the neutrino spin-flavor precession is $P_{\beta\to\bar{\alpha}}\propto|\mathcal{M}_{\beta\to\bar{\alpha}}|^{2}$. In case of the Majorana neutrinos motion in vacuum, the propagators are diagonal, $\Sigma_{ab}\propto\delta_{ab}$. In this situation, the process $\beta\to\bar{\alpha}$ corresponds to neutrino-to-antineutrino oscillations, as we mentioned earlier. The probability of $\nu_{\alpha}\to\bar{\nu}_{\beta}$ oscillations was found in Refs. [14, 24] to be of the order $\sim(m_{1,2}/E)^{2}$.
If Majorana neutrinos have a transition magnetic moment, as one can see in Eq. (2), the wave equations for $\eta_{1}$ and $\eta_{2}$ are coupled by the magnetic interaction. This interaction induces the transitions between $\eta_{1,2}$ and $\mathrm{i}\sigma_{2}\eta_{2,1}^{*}=\xi_{2,1}$. It means that both diagonal, $\Sigma_{11}$ and $\Sigma_{22}$, and nondiagonal, $\Sigma_{12}$ and $\Sigma_{21}$, propagators in Eq. (3.10) can contribute to the transition probability. In case we study the spin-flavor precession like $\nu_{e}\to\bar{\nu}_{\mu}$, the matrix element in Eq. (3.9) takes the form,
$\mathcal{M}_{e\to\bar{\mu}}=\kappa_{\bar{\mu}\mathrm{R}}^{\dagger}(p_{\mu})\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}e^{\mathrm{i}\mathbf{qL}}\left[\frac{1}{2}\left(\Sigma_{22}-\Sigma_{11}\right)\sin 2\theta+\cos^{2}\theta\Sigma_{21}-\sin^{2}\theta\Sigma_{12}\right]\kappa_{e\mathrm{L}}(p_{e}),$ (3.11)
@@ 4 Propagators of massive Majorana neutrinos in a magnetic field
In this section, we derive the dresses propagators of Majorana neutrinos interacting with a magnetic field. For this purpose, we use the method developed in Ref. [20], where the propagators in question are the solutions of the Dyson equations exactly accounting for the interaction with an external field. First, we derive the Dyson equations within the perturbative approach, with all the terms in the perturbation series being summed up. Then, the Dyson equations are solved exactly for ultrarelativistic neutrinos. We deal with two mass eigenstates since we mentioned in Ref. [20] that extension of the formalism for the greater number of neutrinos is ambiguous.
@@
Figure 2: Feynman diagrams corresponding to Eq. (4). 4D Fourier images of the bare propagators $\mathrm{i}S_{a\mathrm{R,L}}$, given in Eqs. (A.2) and (A.10), are depicted by thin solid lines [25, pp. 468–472]. The broad solid lines correspond to the dressed propagators $\mathrm{i}\Sigma_{ab}$. The potential of the magnetic interaction is $V=\mathrm{i}\mu(\bm{\sigma}\mathbf{B})$.
The series in Eq. (4) can be summed up, giving the result
@@
We notice the diagonal propagators in Eq. (4) do not contribute to the matrix element since $\Sigma_{aa}$ preserves the neutrino helicity, whereas $\mathcal{M}_{e\to\bar{\mu}}$ implies the spin-flip. We provide the detailed calculation for $\Sigma_{12}$ and $\Sigma_{21}$ only.
[…]
We provide the detailed calculations of $\Sigma_{12}$ only. Using Eqs. (A.2) and (A.10), the exact solution of the Dyson equation for $\Sigma_{12}$ in Eq. (4) has the form,
@@
$\times\big{\{}\alpha(\varsigma_{1}-\varsigma_{2})\mu(\hat{q}\mathbf{B})+(\bm{\sigma}\cdot[\mathrm{i}\alpha(\varsigma_{1}+\varsigma_{2})\mu(\hat{q}\times\mathbf{B})-2\alpha\varsigma_{1}\varsigma_{2}\mu(\hat{q}\mathbf{B})\hat{q}$
@@
and $E_{1,2}=\sqrt{q^{2}+m_{1,2}^{2}}$. For ultrarelativistic neutrinos, one has that $\varsigma_{1,2}\to 1$, and Eq. (4) takes the form,